# Complex Saddles of Charged-AdS Gravitational partition function

### v1

@@ II.3 Thermodynamic phase-transition and instability
Discontinuous changes to the orientation of Lefschetz thimbles/intersection number are related to the phase transition and thermodynamic instability of the black holes. To justify the statement, let us examine the behaviour of specific heat ($C$), which is given by
@@
The specific heat at the black hole saddles is
@@
At high temperatures ($\beta<\beta_{\rm max}$), the specific heat $C$, of small black holes ($r_{+}^{-}$) is negative, while for large black holes ($r_{+}^{+}$) it is positive. This indicates the large black hole is thermodynamically stable, while the small black hole is unstable. We also observed that the thimbles ($\mathcal{K}(\rm s.b.h)$) emanating from a small black hole exhibit a jump in orientation across the Stokes ray, resulting in a discontinuous change in the intersection number ($n_{\sigma}$). This turns out to be true for the charged case as well, where a negative specific heat saddle corresponds to a jump in $n_{\sigma}$, (see sec. III.1.5 and III.2.4). As a consequence, their contribution to the partition function drops out upon homology averaging. At low temperatures ($\beta>\beta_{\rm max}$), $C$ becomes complex, signifying the solutions describing complex black holes or naked singular geometries. A slightly related conclusion is also mentioned earlier in Marolf:2022ybi .
@@
Let us also point out that the specific heat diverges (dictating continuous phase transition) when $T^{\prime}(r_{+})=0$, which happens when the saddles coalesce at $\beta=\beta_{\rm max}$. As discussed earlier, when one crosses this point, the orientation of the Lefschetz thimbles ($\mathcal{J}$) changes discontinuously, see eq. (37). Hence, we observe that any discontinuous jump in the orientation of thimbles is associated with unphysical behaviour of specific heat.
@@ III Charged AdS Schwarzschild black hole
We next consider the case of charged AdS geometries to test the observations made in the previous section on pure-AdS: naked-singular geometries which though being dominant in partition function in some cases gets eventually discarded in the Picard-Lefschetz analysis of path-integral and small size sub-dominant blackhole geometries despite being relevant drops out from the path-integral due to homology-averaging.
@@
These geometries, which correspond to black hole solutions, consist of inner and outer horizons at $r_{-}$ and $r_{+}$, respectively. These can be found using $f(r)=0$, smaller positive real root of which correspond to inner horizon. As $r_{+}$ satisfies $f(r_{+})=0$, this immediately gives a relation between
@@
Note that the presence of a real and positive $r_{+}$ automatically ensures another real and positive root ($r_{-}$, inner horizon). This is a consequence of Descartes’ rule of signs that constrains $f(r)=0$ to have either two or zero real and positive roots. For a generic black hole geometry, to have a non-singular horizon and avoid the presence of naked singular geometries in the path integral, one requires
@@
to be satisfied by a real positive $r_{+}$. The inequality in eq. (52) also ensures the positivity of $\beta$ at the horizon, see fig (6). We point out that violating eq. 52 will immediately give a naked singular geometry; however, satisfying it doesn’t guarantee the absence of a naked singularity. For instance, a negative saddle or a generic complex saddle can satisfy the condition (52). However, they all correspond to a naked singularity, which is in conflict with the cosmic censorship. We will address this issue more systematically by investigating such naked singular geometries that appear as dominant saddle geometries of the gravitational path integral in the semiclassical limit.
@@
We analyze the partition function for the charged case in two different settings. The first corresponds to fixing the inverse temperature $\beta$ and the electrostatic potential $\Phi$ at spatial infinity, known as the grand canonical ensemble. The other case is where a fixed charge of the spacetime and asymptotically fixes the temperature, known as the canonical ensemble.
@@ III.1 Fixed Potential - Grand Canonical Ensemble
In region $\mathbb{D}$, the integrand has no essential singularity. To start with, we allow all possible charges and take the integration contour for $q$ to run from -$\infty$ to $\infty$, as there is no classical instability present (unlike the rotating solutions Hawking:1999dp ). However, as we explain below at saddles, $q$ and $r_{+}$ get constrained and satisfy eq. (52). The integral in eq. (55) is a two-dimensional integral. However, since the argument inside the exponent is quadratic in $q$ (and hence, there is only one saddle for it), one can first perform the integration over $q$ and retain the leading order term in $G$, reducing to an effective one-dimensional integral, which is appropriate to do for our analysis in the semi-classical limit. The effective one-dimensional integral for $r_{+}$ is then given by:
@@
$\frac{\partial\mathcal{I}}{\partial r_{+}}\bigg|_{r_{+}^{s}}=0\,$, (58)
that gives the relation satisfied by the $r_{+}^{s}$ as:
@@
The saddles are thermal-AdS at $r_{+}^{(0)}=0$ with $q^{(0)}=0$ having a degeneracy of (d3)(d-3), along with a two more saddles denoted $r_{+}^{-}$ and $r_{+}^{+}$. These saddles have the following expression
@@
with associated charge $q^{\pm}=\gamma\Phi(r_{+}^{\pm})^{d-2}$. This can be obtained from the regularity condition on vector potential $A_{\tau}(r_{+})=0$ at the horizon. Defining ($\beta_{\rm max}^{\Phi}$, $\Phi_{\rm max}$)- a special value of external parameters where one observes qualitative changes in the saddles and Lefschetz thimbles
@@
These saddles in eq. (62) share the same structure as the saddles for the AdS Schwarzschild case, given in eq. (15) with the only exception that $\beta_{\rm max}^{\Phi}$ can become imaginary depending on $\Phi$. These saddles are regular for all $\beta$ and $\Phi$. The Kretschmann scalar for these saddle geometries can be computed as before (see appendix A), and it is given by
@@
and $\mu$ is mentioned in eq. (51). It is evident that at $r=0$, the Kretschmann scalar blows up for all saddle geometries. If $\theta$ becomes complex, which happens for real $\beta_{\rm max}^{\Phi}$ ($\Phi<\Phi_{\rm max}$) with $\beta>\beta_{\rm max}^{\Phi}$, the saddles $r_{+}^{\pm}$ become complex. These saddles lack horizon exposing the singularity at $r=0$. These correspond to naked-singular geometries. Real solutions for $r_{+}$ occur under three circumstances: (1) $\Phi>\Phi_{\rm max}$, leading to imaginary $\beta_{\rm max}^{\Phi}$ giving real $\theta$ for values of $\beta$ (2) Extremal case when $\Phi=\Phi_{\rm max}$, implying $\beta_{\rm max}^{\Phi}\to\infty$ leading to two real solutions for all $\beta$ (3) $\Phi<\Phi_{\rm max}$ with $\beta<\beta_{\rm max}^{\Phi}$. In the first case when $\Phi>\Phi_{\rm max}$, one has $r_{+}^{+}>0$ and $r_{+}^{-}<0$. The case where both $r_{+}^{\pm}$ are real and non-negative correspond to second and third case above.
@@
Similar to the AdS Schwarzschild case, the spacetime geometries at the charged saddle points $r_{+}^{\pm}$ are regular, i.e. $\beta=4\pi/f^{\prime}(r_{+})$ is satisfied, implying there is no conical singularity at the horizon. When $r_{+}^{\pm}$ is a positive saddle, utilizing $q=\gamma\Phi(r_{+}^{\pm})^{d-2}$, one satisfies the condition for the no-naked singularity given in eq. (52).
Figure 7: Phase diagram for charged $AdS_{5}$ black hole. The blue shaded region describes a black hole with a non-singular horizon. Region-II has complex saddles corresponding to naked singular geometries.
Similar to the AdS–Schwarzschild case, one encounters a similar puzzle in the grand canonical ensemble as the complex saddles dominate at extremely low temperatures, $\beta\gg\beta_{\rm max}^{\Phi}$, for $\Phi<\Phi_{\rm max}$. The analysis proceeds analogously to the Schwarzschild case, with $\beta_{\rm max}$ replaced by $\beta_{\rm max}^{\Phi}$. Furthermore, for the grand canonical case, in the regime $\Phi>\Phi_{\rm max}$, there exists a negative saddle. At extreme low temperatures $\beta\to\infty$, it remains subdominant with respect to the positive black hole saddle. However, its dominance relative to the thermal AdS varies with the dimension: the negative saddle dominates over the thermal AdS in even $d$, while in odd $d$, the thermal AdS phase remains dominant. Similar to earlier case, we now proceed to analyze these saddle points to determine their contributions to the partition function across different parameter regimes of $\beta$ and $\Phi$, as summarized in the table. (2), utilizing Picard-Lefschetz (PL) methods.
@@ III.1.1 Complex saddles - $\beta>\beta_{\rm max}^{\Phi}$ and $\Phi<\Phi_{\rm max}$
When $\Phi<\Phi_{\rm max}$ and $\beta>\beta_{\rm max}^{\Phi}$ (low temperature), we obtain thermal-AdS along with two complex saddles, similar to the AdS-Schwarzschild case. In a similar argument discussed in sec. II.1, these complex saddles, although sometimes dominant, do not contribute to the partition function, according to the PL method, see fig 1. It can be quickly understood as follows: in the low-temperature regime, the complex saddle points yield a nonzero value of $\mathcal{H}(r_{+}^{\pm},\Phi^{\pm})$ and fails to satisfy the relevance condition $\mathcal{H}(r_{+}^{\pm},\Phi^{\pm})=0$ in any dimension $d\geq 3$. Only the thermal saddle contributes as it lies on the integration contour. This implies that the complex saddles corresponding to naked-singular geometries, despite being dominant, fail to become “relevant” saddles as the deformed integration contour, according to the PL method, doesn’t pass through them.
@@ III.1.2 Negative real saddle - $\Phi>\Phi_{\rm max}$
When $\Phi>\Phi_{\rm max}$, at any temperature, we have thermal AdS, a real positive $(r_{+}^{+})$ and a negative saddle $(r_{+}^{-})$. The positive saddle corresponds to a black hole with a horizon. However, the negative saddle corresponds to a naked singularity. Since all the saddles lie on the real axis, we have $\mathcal{H}(r_{+}^{s})=0$ for each of them and hence saddles lie on a Stokes ray. In a way similar to the AdS-Schwarzschild case, we complexify the Newton constant ($G=|G|e^{i\epsilon}$) and perform the PL analysis (see fig. 9). In the regime $\Phi>\Phi_{\rm max}$, we find that both the thermal-AdS saddle and the positive black hole saddle are relevant for either sign of $\epsilon$, while the negative saddle doesn’t contribute since the associated steepest ascent doesn’t intersect the original integration contour. Hence, we reach the same conclusion that naked singular geometries do not contribute to the partition function.
@@ III.1.4 Extremal case - $\Phi=\Phi_{\rm max}$
For the extremal case, one finds the saddle solutions are the thermal AdS saddle (now with the degeneracy factor of d2d-2, as one of the black hole saddles coincides with thermal ads) and a black hole saddle with horizon at $r_{+}=4\pi l^{2}/(d\beta)$ which exists for all temperatures ($0<\beta<\infty$). Since saddles are real, they lie on the Stokes ray, and as before, one can complexify $G$ and proceed with PL analysis (see fig. 10).
@@ III.1.5 Thermodynamic instability and homology jump
Similar to the AdS Schwarzschild, we examine the specific heat ($C_{\Phi}$) of a charged blackhole at constant potential ($\Phi$) and susceptibility ($\chi_{\beta}$) and analyze how the unphysical behaviour of them is related to the discontinuous change in the orientation of the thimbles/intersection number. The specific heat of the black hole at constant potential ($\Phi$) is given by:
@@
In the regime $\beta<\beta_{\rm max}^{\Phi}$ and $\Phi<\Phi_{\rm max}$, $C_{\Phi}$ is positive for large black hole ($r_{+}^{+}$) and negative for small black hole $(r_{+}^{-})$. Hence, a large black hole is thermodynamically stable while a small black hole is unstable. At low temperatures $\beta>\beta_{\rm max}^{\Phi}$ and $\Phi<\Phi_{\rm max}$, $C_{\Phi}$ becomes complex, signifying that the corresponding solutions describe complex black holes or naked singular geometries, which are unphysical. This scenario has been demonstrated in the fig. (11(a)). The specific heat $(C_{\Phi})$ diverges, when $\beta=\beta_{\rm max}^{\Phi}$. For the small black hole branch, it diverges to -$\infty$ while for the large black hole branch, it diverges to $\infty$. As discussed earlier, as one crosses the critical point $\beta=\beta_{\rm max}^{\Phi}$, the orientation of the Lefschetz thimbles changes discontinuously.
Similar to the AdS-Schwarzchild, charged small black holes are always subdominant than ads and the large black hole, and hence always suffer from the Stokes jump. The Lefschetz thimbles associated with small black holes undergo a homology jump across the Stokes ray. Hence, upon taking the homology average, their contribution is eliminated from the partition function. Similarly, in the regime $\Phi>\Phi_{\rm max}$, the specific heat $C_{\Phi}$ associated with the positive black hole saddle remains positive, indicating thermodynamic stability. In contrast, the negative black hole saddle exhibits a negative specific heat for odd $d$ and a positive specific heat for even $d$. When $d$ is odd, the negative saddle is subdominant to both the ads and the positive saddle and hence shows a similar homology jump in the Lefschetz thimbles. As discussed earlier, the negative saddle corresponds to a naked singular geometry and, according to the Picard–Lefschetz technique, does not contribute to the partition function.
@@
$\chi_{\beta}(r_{+})=\frac{\omega(d-2)r_{+}^{d-3}}{4\pi G}\frac{d(d-3)\beta r_{+}^{2}+(2d-5)\pi l^{2}r_{+}+\beta l^{2}(d-2)^{2}}{d\beta r_{+}-2\pi l^{2}}\,.$ (70)
It is easy to see from the above equation for $d\geq 3$, at high temperatures, the susceptibility for the small black hole $r_{+}^{-}<2\pi l^{2}/(d\beta)$ is negative, while large black holes $r_{+}^{+}>2\pi l^{2}/(d\beta)$ is positive. This indicates again that the small black hole is thermodynamically unstable, while the large black hole is stable. The susceptibility diverges when both black holes coalesce to one with horizon radius $r_{+}=2\pi l^{2}/d\beta$, which corresponds to $\beta=\beta_{\rm max}^{\Phi}$ where $C_{\Phi}$ also diverges. This situation is illustrated in the figure 11(b). While at low temperatures, the susceptibility at the complex saddle solution is complex. However, the susceptibility of negative solution alter sign as we go through different dimensions. To see this explicitly, utilizing the saddle equation for negative saddle $r_{+}=-|r_{+}|$, one can recast the $\chi_{\beta}$ in eq. (70) as:

### v2

@@ II.3 Thermodynamic phase-transition and instability
In this section, we wish to understand the thermodynamic phase transitions and instabilities from the viewpoint of Picard-Lefschetz analysis Witten:2010cx ; Kanazawa:2014qma ; Fujimori:2021oqg . We observe that any discontinuous changes to the orientation of Lefschetz thimbles/intersection number are related to the phase transition and thermodynamic instability of the black holes. To justify the statement, let us examine the behaviour of specific heat ($C$), which is given by
@@
The specific heat of the black hole saddles given in eq. (15) is
@@
At high temperatures ($\beta<\beta_{\rm max}$), one finds the specific heat $C$, of small black holes ($r_{+}^{-}$) is negative, while for large black holes ($r_{+}^{+}$) it is positive, see fig. (5). This indicates the large black hole is thermodynamically stable, while the small black hole is unstable. We also observed that the thimbles ($\mathcal{K}(\rm s.b.h)$) emanating from a small black hole exhibit a jump in orientation across the Stokes ray, resulting in a discontinuous change in the intersection number ($n_{\sigma}$). This turns out to be true for the more general charged case as well, where a negative specific heat saddle corresponds to a jump in $n_{\sigma}$, (see sec. III.1.5 and III.2.4). As a consequence, their contribution to the partition function drops out upon homology averaging. At low temperatures ($\beta>\beta_{\rm max}$), $C$ becomes complex, signifying the solutions describing complex black holes or naked singular geometries. A similar conclusion is also observed earlier in Marolf:2022ybi , where the authors also noticed that when black hole saddles have positive specific heat, they contribute to the partition function with non-zero weight in the semi-classical limit.
@@
Let us also point out that the specific heat diverges (dictating the continuous phase transition) when $T^{\prime}(r_{+})=0$, which happens when the saddles coalesce at $\beta=\beta_{\rm max}$, see fig. (5). As discussed earlier, when one crosses this point, the orientation of the Lefschetz thimbles ($\mathcal{J}$) changes discontinuously, see eq. (37). Hence, we find that any discontinuous jump in the orientation of thimbles is associated with unphysical behaviour of specific heat and leads to thermodynamic instability.
@@ III Charged AdS Schwarzschild black hole
We next consider the case of charged AdS geometries to test the observations made in the previous section on pure-AdS: naked-singular geometries, although being dominant in partition function in some cases, are eventually discarded in the Picard-Lefschetz analysis and small size sub-dominant blackhole geometries despite being relevant, drops out from the partition function due to homology-averaging.
@@
The geometries that correspond to black hole solutions, consist of inner and outer horizons at $r_{-}$ and $r_{+}$, respectively. These can be found using $f(r)=0$, where smaller positive real root correspond to inner horizon and larger positive real root correspond to outer horizon. As $r_{+}$ satisfies $f(r_{+})=0$, this immediately gives a relation between $\mu$ and $q$
@@
Note that the presence of a real and positive $r_{+}$ automatically ensures another real and positive root ($r_{-}$, inner horizon). This is a consequence of Descartes’ rule of signs that constrains $f(r)=0$ to have either two or zero real and positive roots. For a generic black hole geometry, to have a non-singular horizon and avoid the presence of naked singular geometries in the partition function, one requires
@@
to be satisfied by a real positive $r_{+}$. The inequality in eq. (52) saturates when both the horizons coincide and one gets extremal black hole, see fig (6). The inequality ensures the positivity of $\beta$ at the horizon ($r_{+}$) and also put a bound on the mass ($\mu\geq\mu_{e}$), where $\mu_{e}$ is the mass of extremal black hole. In supersymmetric theories, one also gets a bound on the mass which is $\mu\geq 2q$, with $\mu=2q$ is a BPS state London:1995ib ; Chamblin:1999tk . Now, since $\mu_{e}>2q$ for finite $l$, extremal solution is always non-supersymmetric. When $\mu=2q$ (supersymmetric solution), one gets from eq. (51)
$\left(1+\frac{q}{r_{+}^{d-2}}\right)^{2}+\frac{r_{+}^{2}}{l^{2}}=0.$ (53)
Clearly, eq. (53) has no real $r_{+}$ solutions for real $q$ and hence will describe a naked singularity Chamblin:1999tk . Such solutions will also violate the inequality, 52. We point out that all solutions violating eq. 52 will correspond to a naked singular geometry; however, satisfying it doesn’t always guarantee the absence of a naked singularity. For instance, real negative saddles will satisfy the condition (52). However, they all correspond to a naked singularity, which is in conflict with the cosmic censorship. We will address this issue more systematically by investigating the relevance of such geometries that appear as dominant saddle geometries of the gravitational partition function in the semiclassical limit.
@@
We analyze the partition function for the charged case in two different settings. The first corresponds to fixing the inverse temperature $\beta$ and the electrostatic potential $\Phi$ at spatial infinity, known as the grand canonical ensemble. The other case is where a fixed charge ($Q)$ and $\beta$ asymptotically, known as the canonical ensemble.
@@ III.1 Fixed Potential - Grand Canonical Ensemble
In region $\mathbb{D}$, the integrand has no essential singularity. To start with, we allow all possible charges and take the integration contour for $q$ to run from -$\infty$ to $\infty$, as there is no classical instability present (unlike the rotating solutions Hawking:1999dp ). However, as we explain below at saddles, $q$ and $r_{+}$ get constrained and satisfy eq. (52). The integral in eq. (56) is a two-dimensional integral. Since the argument inside the exponent is quadratic in $q$ (and hence, there is only one saddle for it), one can first perform the integration over $q$ and retain the leading order term in $G$, reducing to an effective one-dimensional integral. This is appropriate to do for our analysis in the semi-classical limit. The effective one-dimensional integral for $r_{+}$ is then given by:
@@
$\frac{\partial\mathcal{I}}{\partial r_{+}}=0\,$, (59)
that gives the relation satisfied by the $r_{+}$ as:
@@
The saddles are thermal-AdS at $r_{+}^{(0)}=0$ with $q^{(0)}=0$ having a degeneracy of (d3)(d-3), along with two more saddles denoted $r_{+}^{-}$ and $r_{+}^{+}$. These saddles have the following expression
@@
with associated charge $q^{\pm}=\gamma\Phi(r_{+}^{\pm})^{d-2}$. This is obtained from the regularity condition on vector potential $A_{\tau}(r_{+})=0$ at the horizon. Defining ($\beta_{\rm max}^{\Phi}$, $\Phi_{\rm max}$)- a special value of external parameters where one observes qualitative changes in the saddles and Lefschetz thimbles
@@
The saddles in eq. (63) share the same structure as the saddles for the AdS Schwarzschild case, given in eq. (15) with the only exception that $\beta_{\rm max}^{\Phi}$ can become imaginary depending on $\Phi$. These saddles are regular for all $\beta$ and $\Phi$. The Kretschmann scalar for these saddle geometries can be computed as before (see appendix A), and it is given by
@@
and $\mu$ is mentioned in eq. (51). It is evident that at $r=0$, the Kretschmann scalar blows up for all saddle geometries. If $\theta$ becomes complex, which happens for real $\beta_{\rm max}^{\Phi}$ ($\Phi<\Phi_{\rm max}$) with $\beta>\beta_{\rm max}^{\Phi}$, the saddles $r_{+}^{\pm}$ become complex. These saddles lack a horizon, exposing the singularity at $r=0$. These correspond to naked-singular geometries. Real solutions for $r_{+}$ occur under three circumstances: (1) $\Phi>\Phi_{\rm max}$, leading to imaginary $\beta_{\rm max}^{\Phi}$ giving real $\theta$ for values of $\beta$, (2) Extremal case when $\Phi=\Phi_{\rm max}$, implying $\beta_{\rm max}^{\Phi}\to\infty$ leading to two real solutions for all $\beta$, (3) $\Phi<\Phi_{\rm max}$ with $\beta<\beta_{\rm max}^{\Phi}$. In the first case when $\Phi>\Phi_{\rm max}$, one has $r_{+}^{+}>0$ and $r_{+}^{-}<0$. The case where both $r_{+}^{\pm}$ are real and non-negative correspond to second and third case above.
@@
Similar to the AdS Schwarzschild case, the spacetime geometries at the charged saddle points $r_{+}^{\pm}$ are regular, i.e. $\beta=4\pi/f^{\prime}(r_{+})$ is satisfied, implying there is no conical singularity at the horizon. Further, when $r_{+}^{\pm}$ is a positive saddle, utilizing $q=\gamma\Phi(r_{+}^{\pm})^{d-2}$, one satisfies the condition for the no-naked singularity given in eq. (52).
Figure 7: Phase diagram for charged $AdS_{5}$ black hole. The blue shaded region describes naked singular geometries. Region-II describes a black hole with a non-singular horizon.
Just like in the case of AdS–Schwarzschild, one encounters a similar puzzle in the grand canonical ensemble that is the complex saddles dominate at extremely low temperatures, $\beta\gg\beta_{\rm max}^{\Phi}$, for $\Phi<\Phi_{\rm max}$. The analysis proceeds analogously to the Schwarzschild case, with $\beta_{\rm max}$ replaced by $\beta_{\rm max}^{\Phi}$. Furthermore, for the grand canonical case, in the regime $\Phi>\Phi_{\rm max}$, there exists a negative saddle. At extreme low temperatures $\beta\to\infty$, it remains subdominant with respect to the positive black hole saddle. However, its dominance relative to the thermal AdS varies with the dimension: the negative saddle dominates over the thermal AdS in even $d$, while in odd $d$, the thermal AdS phase remains dominant. Similar to earlier case, we now proceed to analyze these saddle points to determine their contributions to the partition function across different parameter regimes of $\beta$ and $\Phi$, as summarized in the table. (2), utilizing Picard-Lefschetz (PL) methods.
@@ III.1.1 Complex saddles - $\beta>\beta_{\rm max}^{\Phi}$ and $\Phi<\Phi_{\rm max}$
When $\Phi<\Phi_{\rm max}$ and $\beta>\beta_{\rm max}^{\Phi}$ (low temperature), we obtain thermal-AdS along with two complex saddles, similar to the AdS-Schwarzschild case. In a similar argument discussed in sec. II.1, these complex saddles, although sometimes dominant, do not contribute to the partition function, according to the PL method, see fig 2. It can be quickly understood as follows: in the low-temperature regime, the complex saddle points yield a nonzero value of $\mathcal{H}(r_{+}^{\pm},\Phi)$ and fails to satisfy the relevance condition $\mathcal{H}(r_{+}^{\pm},\Phi)=0$ in any dimension $d\geq 3$. Only the thermal saddle contributes as it lies on the integration contour. This implies that the complex saddles corresponding to naked-singular geometries, despite being dominant, fail to become “relevant” saddles as the deformed integration contour, according to the PL method, doesn’t pass through them.
@@ III.1.2 Negative real saddle - $\Phi>\Phi_{\rm max}$
When $\Phi>\Phi_{\rm max}$, at any temperature, we have thermal AdS, a real positive $(r_{+}^{+})$ and a negative saddle $(r_{+}^{-})$. The positive saddle corresponds to a black hole with a horizon. However, the negative saddle corresponds to a naked singularity. Since all the saddles lie on the real axis, we have $\mathcal{H}(r_{+}^{s})=0$ for each of them and hence saddles lie on a Stokes ray. In a way similar to the AdS-Schwarzschild case, we complexify the Newton constant ($G=|G|e^{i\epsilon}$) and perform the PL analysis (see fig. 9). In this regime $\Phi>\Phi_{\rm max}$, we find that both the thermal-AdS saddle and the positive black hole saddle are relevant for either sign of $\epsilon$, while the negative saddle doesn’t contribute since the associated steepest ascent doesn’t intersect the original integration contour. Hence, we reach the same conclusion that naked singular geometries do not contribute to the partition function.
@@ III.1.4 Extremal case - $\Phi=\Phi_{\rm max}$
For the extremal case, one finds the saddle solutions are the thermal AdS saddle (now with the degeneracy factor of d2d-2, as one of the black hole saddles coincides with thermal ads) and a black hole saddle with horizon at $r_{+}=4\pi l^{2}/(d\beta)$ which exists for all temperatures ($0<\beta<\infty$). Since saddles are real, they lie on the Stokes ray, and as before, one can complexify $G$ and proceed with PL analysis (see fig. 10). We find that both the saddles contribute to the partition function for either sign of $G$ - rotation.
@@ III.1.5 Thermodynamic instability and homology jump
In this section, we examine the specific heat ($C_{\Phi}$) at constant potential ($\Phi$) and susceptibility ($\chi_{\beta}$) of a charged black hole, extending the analysis done in sec. (II.3) for the Schwarzschild black hole. We wish to understand how the unphysical behaviour of these thermodynamic quantities is related to the discontinuous change in the orientation of the thimbles/intersection number. The specific heat of the black hole at constant potential ($\Phi$) is given by:
@@
In the regime $\beta<\beta_{\rm max}^{\Phi}$ and $\Phi<\Phi_{\rm max}$, $C_{\Phi}$ is positive for large black hole ($r_{+}^{+}$) and negative for small black hole $(r_{+}^{-})$. Hence, the large black hole is thermodynamically stable, while the small black hole is unstable. At low temperatures $\beta>\beta_{\rm max}^{\Phi}$ and $\Phi<\Phi_{\rm max}$, $C_{\Phi}$ becomes complex, signifying that the corresponding solutions describe complex black holes or naked singular geometries, which are unphysical. This scenario has been demonstrated in the fig. (11(a)). The specific heat $(C_{\Phi})$ diverges, when $\beta=\beta_{\rm max}^{\Phi}$. For the small black hole branch, it diverges to -$\infty$ while for the large black hole branch, it diverges to $\infty$. As discussed earlier, as one crosses the critical point $\beta=\beta_{\rm max}^{\Phi}$, the orientation of the Lefschetz thimbles changes discontinuously.
Similar to the AdS-Schwarzchild, charged small black holes are always subdominant than AdS and the large black hole, and hence always suffer from the Stokes jump. The Lefschetz thimbles associated with small black holes undergo a homology jump across the Stokes ray. Upon taking the homology average, their contribution is eliminated from the partition function. Similarly, in the regime $\Phi>\Phi_{\rm max}$, the specific heat $C_{\Phi}$ associated with the positive black hole saddle remains positive, indicating thermodynamic stability. In contrast, the negative black hole saddle exhibits a negative specific heat for odd $d$ and a positive specific heat for even $d$. When $d$ is odd, the negative saddle is subdominant with respect to both the AdS and the positive saddle and therefore exhibits a similar homology jump in the Lefschetz thimbles. As discussed earlier, the negative saddle corresponds to a naked singular geometry and, according to the Picard–Lefschetz technique, does not contribute to the partition function.
@@
$\chi_{\beta}(r_{+})=\frac{\omega(d-2)r_{+}^{d-3}}{4\pi G}\frac{d(d-1)\beta r_{+}^{2}-2(2d-3)\pi l^{2}r_{+}+\beta l^{2}(d-2)^{2}}{d\beta r_{+}-2\pi l^{2}}\,.$ (71)
It is easy to see from the above equation for $d\geq 3$, at high temperatures, the susceptibility for the small black hole $r_{+}^{-}<2\pi l^{2}/(d\beta)$ is negative, while large black holes $r_{+}^{+}>2\pi l^{2}/(d\beta)$ is positive. This indicates again that the small black hole is thermodynamically unstable, while the large black hole is stable. The susceptibility diverges when both black holes coalesce to one with horizon radius $r_{+}=2\pi l^{2}/d\beta$, which corresponds to $\beta=\beta_{\rm max}^{\Phi}$ where $C_{\Phi}$ also diverges. This situation is illustrated in the figure 11(b). At low temperatures, the susceptibility at the complex saddle solution is complex. However, the susceptibility of negative solution alter sign as we go through different dimensions. To see this explicitly, utilizing the saddle equation for negative saddle $r_{+}=-|r_{+}|$, and recast the $\chi_{\beta}$ in eq. (71) as: